) A gun barrel weight 300 kg and has a recoil spring of

stiffness 250N/mm. The barrel recoils 0.8m on firing.

Determine the (i) Critical recoil velocity of the gun,

(ii) Critical damping coefficient of the dashpot engaged

at the end of the recoil stroke.


1
Expert's answer
2021-02-05T04:56:06-0500

natural frequency

πœ” = "\sqrt \frac{𝑘}{𝑚}"

= "\sqrt \frac{250000}{300}"

= 28.867π‘Ÿπ‘Žπ‘‘/𝑠 .


From conservation of energy,

"\frac{𝑚𝑣2}{2} = \frac{kx2}{2}"

𝑣 = πœ”π‘₯

= (28.867)(0.8)

= 23.09396 π‘š/ 𝑠 .